OpenStudy (anonymous):

Suppose y varies directly with x. Write an equation relating x and y. y=12 when x=4 ..

OpenStudy (anonymous):

When y directly varies with x then it simply means that: \[y = k \cdot x\] Where, \(k\) is the proportionality constant..

OpenStudy (anonymous):

Now, find k when you are given with x = 4 and y = 12.. Can you find k here??

OpenStudy (anonymous):

wait what ? o.O

OpenStudy (anonymous):

So, you have: \(y = k \cdot x\) Now plug in here y = 12 and x = 4, and find \(k\) here.

OpenStudy (anonymous):

a. y = 3x c. y = 4x b. y = x d. y = − 3x so where does 3 come in on these answers ?

OpenStudy (anonymous):

Find \(k\) there and you will know 3 came there.

OpenStudy (anonymous):

*how.

OpenStudy (anonymous):

how do you find k tho? im sorry im a slow learner .....

OpenStudy (anonymous):

Okay let us go slow then.. Did you get this: \[y = kx\] ??

OpenStudy (anonymous):

no ..

OpenStudy (anonymous):

When we say two quantities are directly proportional or like x varies directly with y, then it means: \[x = k \times y\] But in our case it is given that , y varies directly with x, then you will have: \[y = k \times x\] Now getting?

OpenStudy (anonymous):

kind of

OpenStudy (anonymous):

Make it sure first.

OpenStudy (anonymous):

i really dont get it i dont get anything and my time for my test is almost up ima just guess thnks for trying :) ..

OpenStudy (anonymous):

Giving test?? Oh sorry, then I can't help you more.. But when you will actually want to learn, then do mention me in your posts.. You are welcome dear..

OpenStudy (anonymous):

oh trust me i wont ha

OpenStudy (anonymous):

Good for you.. :)